Convert the following two equations to slope-intercept form:

a) 1/3x + 3 = 3/4y
b) 3x -4y - 20 = 0

Answer by nyc_function(2741)

TIP: Post one question at a time or else no one will reply.
3x -4y - 20 = 0
3x - 4y = 20
-4y = -3x + 20
y = (-3x + 20)/-4
y = (3/4)x - 5
Done!